i was seven times as old as you were when i was as old as you are said a dad to his son....the sum of thr present ages is 110.....find present age of dad and son
Answer by ankor@dixie-net.com(22740)

i was seven times as old as you were, when i was as old as you are.
said a dad to his son....the sum of their present ages is 110.....find present age of dad and son
:
Let f = father's present age
Let s = son's present age
Let d = (f-s), the difference in their ages
:
"i was seven times as old as you were when i was as old as you are"
f - d = 7(s - d)
f - d = 7s - 7d
f = 7s - 7d + d
f = 7s - 6d
Replace d with (f-s)
f = 7s - 6(f-s)
f = 7s - 6f + 6s
f + 6f = 7s + 6s
7f = 13s
:
"the sum of their present ages is 110."
f + s = 110
f = (110-s)
Replace f with (110-s) in the eq: 7f = 13s
7(110-s) = 13s
770-7s = 13s
770 = 13s + 7s
770 = 20s
s = 770/20
s = 38.5 is the son's present age, (I was hoping for an integer, but anyway:)
then
110-38.5 = 71.5 is Father's age
:
:
See if this checks out in the statement:
71.5 - 38.5 = 33 yr age difference
""i was seven times as old as you were when i was as old as you are"
71.5 - 33 = 7(38.5-33)
38.5 = 7(5.5), seem to check out
:
Father is 71.5, son is 38.5